OpenStudy (anonymous):

height of a tower is 3 times the height of a certain building. if the tower is 50 m taller than the bulding. how tall is the power??

OpenStudy (anonymous):

tower*

OpenStudy (gw2011):

Let B=Building Let T=Tower T=3B and T=B+50 3B=B+50 3B-B=50 2B=50 B=25m (Thisis the height of the Building) T=25+50=75m (This is the height of the Tower)